OpenStudy (anonymous):

Algebra 1 help needed Next, graph y = x3 – 4x2 + 6. Describe what you see.

OpenStudy (anonymous):

Will fan and medal 😋

OpenStudy (danjs):

Do you just need to calculate a few points? If x=0, then y=0-0+6 = 6 First point (x,y) = (0,6)

OpenStudy (danjs):

Let x=1,2,3,-1,-2,-3 Calculate all the Y values, plot the points

OpenStudy (anonymous):

What? I didn't understand what u said

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dx3-4x2%2B6